OpenStudy (anonymous):

There were 170 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $490. How many Juniors bought tickets?

OpenStudy (anonymous):

@godgavemeyou

OpenStudy (anonymous):

I thought I answered this question for you ?

OpenStudy (anonymous):

no u didnt but u can now if u want :)

OpenStudy (anonymous):

hmmm? Let x be the number of Juniors then there are (170-x)number of Senior.... now, 2x+ 3.50(170-x) = 490 solve tht and get the value of x

OpenStudy (anonymous):

how do i solve it?

OpenStudy (anonymous):

cuz i got 90 is that right?

OpenStudy (anonymous):

no its not right

OpenStudy (anonymous):

2x+ 3.50(170-x) = 490 2x+595-3.5x = 490 -1.5x=-105 x=70 What do think the answer is now?

OpenStudy (anonymous):

the answer is 70 right

OpenStudy (anonymous):

x = 70 that is correct !!! Great job don't forget to mark the best response so can show ur thanks!

OpenStudy (anonymous):

i have another question but ill make a new one so u can get a medal :)

OpenStudy (anonymous):

j = juniors s = seniors j + s = 170 --> s = 170 - j 2j + 3.5(170 - j) = 490 2j + 595 - 3.5j = 490 2j - 3.5j = 490 - 595 - 1.5j = - 105 j = -105/-1.5 j = 70 I agree with godgavemeyou...:)

OpenStudy (anonymous):

thanks!! your both great

OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 70 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 240 miles apart?

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

I am not good at these....I am sorry :(

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

maybe godgavemeyou can help

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

im thinking

OpenStudy (anonymous):

i made a new one go to it so i can give u a mtetal

OpenStudy (anonymous):

ok